# MENTOR: Guiding Hierarchical Reinforcement Learning with Human Feedback and Dynamic Distance Constraint

### v1

@@ 5.5 Ablation Studies (RQ4)
In addition to DDC and EED, we employ various techniques such as near-policy sampling in RLHF, and negative sampling in distance model learning and explore encouragement in high-level policy. These ablation studies are presented in Section D.3.
@@ 6 Conclusion
However, we recognize the complexity of human guidance beyond just subgoal selection and aim to explore a wider range of feedback integration to enhance learning dynamics. We plan to further expand the framework’s applications and refine its mechanisms, aiming to advance hierarchical reinforcement learning and create more intuitive, adaptable learning systems.
Appendix A MENTOR details
A.1 execute process
In this section, we will comprehensively explain our framework MENTOR using pseudo-code in Algorithm 1. During the interaction with the environment (from line 5 to line 18), a high-level policy, referred to as $\pi_{hi}$, is responsible for selecting a subgoal $sg$ (line 7). Once the subgoal is determined, a low-level explore policy, denoted as $\pi_{lo}^{e}$, is utilized to execute actions until the subgoal is successfully achieved. This process of selecting subgoals and executing actions continues until the end of the episode. The data obtained from the interaction with the environment, both at the high-level and low-level, is stored in two replay buffers known as $B_{hi}$ and $B_{lo}$. The model update process is implemented in lines 20 to 24. The high-level policy in Algorithm 3 updates $\pi_{hi}$ and $\alpha$ alternately, following equations 9 and 10. As for the low-level policy, it involves updating the RND model, applying hindsight, updating the low-level base policy $\pi_{lo}$, adding the RND bonus in the batch data, and updating the explore policy $\pi_{lo}^{e}$. From line 22 to 23, preference tuple pairs are sampled from the data of the last few episodes in $B_{hi}$ and human labels are obtained and stored in $B_{hf}$. These batches are then used to update $r_{hf}$ and rewrite the reward data in the high-level buffer $B_{hi}$. Finally, distance training data is sampled from the data of the last few episodes in $B_{hi}$ and used to update the distance model $d$. From line 27 to 35, it tests the low-level base policy success rate on the subgoal $sg$ given by high-level policy and adjusts the hyper-parameters $k$.
$\theta_{hi}=\theta_{hi}+\eta_{\pi}\mathbb{E}_{\theta}[\nabla_{\theta}\log\pi(s,sg,g)(\sum r_{hi}(s,sg,g)-\alpha H(sg,g,k))]\big{|}_{\begin{subarray}{c}\theta=\theta_{hi}\\ \pi=\pi_{hi}\end{subarray}}$ (9)
$\alpha=\alpha-\eta_{\alpha}\nabla_{\alpha}\mathbb{E}[-\alpha(d(sg,g)-k)]\big{|}_{\alpha=\alpha}$ (10)
Algorithm 1 MENTOR: High-level Adaptive Control with Dynamic Distance Constraints
1:  Input: initial state distribution $\rho(s)$, goal distribution $\rho(g)$
2:  Initialize high-level policy $\pi_{hi}$, low-level base policy $\pi_{lo}$, low-level explore policy $\pi_{lo}^{e}$, reward model $r_{hf}$, RND model $f_{rnd}$, distance model $d$, balancing coefficient $\alpha$, constraint hyper-parameters $k$, high-level buffer $B_{hi}$, preference buffer $B_{hf}$ and low-level buffer $B_{lo}$.
3:
4:  for $i=0$ to $M$ do
5:     receive initial state $s_{t}$ and initial $g$ from the environment, $s_{t}\leftarrow\rho(s),g\leftarrow\rho(g)$
6:     repeat
7:        select a subgoal $sg$leftarrow $\pi_{hi}(s_{t},g)$
8:        store high-level state $s_{hi}\leftarrow s_{t}$
9:        repeat
10:           select action $a_{t}\leftarrow$ $\pi_{lo}^{e}(s_{t},sg)$
11:           execute $a_{t}$ and observe next state $s_{t+1}$
12:           compute low-level reward $r_{lo}=r_{g}(s_{t+1},sg)$
13:           store $(s_{t},a_{t},s_{t+1},sg,r_{l})$ to low-level buffer $B_{lo}$
14:           $s_{t}\leftarrow s_{t+1}$
15:        until $terminal(s_{t},sg)$
16:        calculate the high reward $r_{hi}\leftarrow r_{hf}(s_{hi},g,sg)$
17:        append $(s_{hi},sg,s_{t},g,r_{hi})$ to $B_{hi}$
18:        $s_{hi}\leftarrow s_{t}$
19:     until end of episode
20:     TrainHighPolicy($\pi_{hi}$, $f_{rnd}$, $d$, $\alpha$, $B_{hi}$) Algorithm 2
21:     TrainLowPolicy($\pi_{lo}$, $\pi_{lo}^{e}$, $f_{rnd}$, $B_{lo}$) Algorithm 3
22:     near policy query preference tuples pairs from $B_{hi}$ and store preferences in $B_{hf}$
23:     update $r_{hf}$ on batches from $B_{hf}$ and rewrite the reward in $B_{hi}$
24:     near policy sample distance batches from $B_{lo}$ and update distance model $d$ using Equation 4
25:     $success\_rate\leftarrow\{\}$
26:     for $j=0$ to $N_{1}$ do
27:        sample goal $g$ from environment and select a subgoal $sg$
[…]
29:     end for
30:     if average($success\_rate$) $\geq$ $high\_threshold$ then
31:        increase $k$ value
32:     end if
33:     if average($success\_rate$) $<$ $low\_threshold$ then
34:        decrease $k$ value
35:     end if
36:  end for
Algorithm 2 TrainHighPolicy
1:  Input: high-level base policy $\pi_{hi}$, distance model $d$, balancing coefficient $\alpha$, high-level buffer $B_{hi}$
2:
3:  for $i=0$ to $N_{2}$ do
4:     sample a batch $B$ from buffer $B_{lo}$
5:     use batch $B$ to update policy $\pi_{hi}$ using Equation 9
6:     use batch $B$ to dual variable $\alpha$ using Equation 10
7:  end for
Algorithm 3 TrainLowPolicy
1:  Input: low-level base policy $\pi_{lo}$, low-level explore policy $\pi_{lo}^{e}$, RND model $f_{rnd}$, low-level buffer $B_{lo}$
2:
3:  for $i=0$ to $N_{3}$ do
4:     sample a batch $B$ from buffer $B_{lo}$
5:     use batch $B$ to update RND model $f_{rnd}$
6:     using hindsight to rewrite the goal and reward using $r_{g}$ in $B$
7:     use batch $B$ to update policy $\pi_{lo}$
8:     add RND bonus $f_{rnd}$ to reward in $B$
9:     use batch $B$ to update explore policy $\pi_{lo}^{e}$
10:  end for
Appendix B Domains Details
@@
$\text{r}_{\text{Four rooms}}(s)=\left\{\begin{array}[]{ll}-\lVert s-[0.25,0.25]\rVert_{2}&\text{if }s_{x}\geq 0\text{ and }s_{y}\geq 0\\ -\lVert s-[0,0.3]\rVert_{2}-0.3&\text{if }s_{x}\leq 0\text{ and }s_{y}\geq 0\\ -\lVert s-[-0.3,0]\rVert_{2}-0.6&\text{if }s_{x}\leq 0\text{ and }s_{y}\leq 0\\ -\lVert s-[0,-0.3]\rVert_{2}-1&\text{if }s_{x}\geq 0\text{ and }s_{y}\leq 0\end{array}\right.$ (11)
Appendix C Implementation Details
In this section, we will present the implementation details of each baseline used in our experiments. We will also discuss the reasonable and necessary modifications made to the original algorithm, as well as provide the rationale behind these modifications.
C.1 MENTOR
The executing process details of MENTOR is introduced in Appendix A. Here, we only provide the hyperparameters.
Table 2: Hyperparameters of MENTOR.
@@ C.1 MENTOR
C.2 HhP
HhP introduces a HRL method that integrates human preferences and curiosity-driven exploration. By using human preferences to guide decision-making at high-levels and curiosity to promote environmental exploration at low-levels. HhP is considered to be inefficient in combining HRL and RLHF, and it also introduces bias at the low-levels. When we trained HhP, we found that this algorithm was difficult to converge, and to ensure the effectiveness of training, we introduced the hindsight technique in this algorithm as well.
C.3 HER
HER enables efficient learning from limited and non-diverse rewards by re-framing failed attempts as successes towards different goals. In our implementation, we utilize a 3-layer MLP policy to incorporate HER. When sampling batch transitions from the replay buffer, we employ hindsight technology to modify certain parts of the transitions. It is important to mention that the high-level policy differs from the low-level policy of MENTOR in certain aspects. Unlike the low-level policy, the high-level policy solely receives goals from the environment and does not incorporate an RND bonus.
C.4 DDL
DDL (dynamic distance learning) calculates dynamical distances, defining them as the expected number of steps between states in a system. This method leverages unsupervised interactions for learning distances. In the distance evaluation step, the aim is to estimate the dynamical distance between pairs of states visited by a given policy. A distance function, denoted as $d(s_{i},s_{j})$, is utilized for this purpose. To calculate the distance, multiple trajectories, denoted as $\tau$, are generated by rolling out the policy. Each trajectory has a length of T. The empirical distance between states $s_{i}$ and $s_{j}$ in the trajectory $\tau$ is then calculated, where $0\leq i\leq j\leq T$. The empirical distance is simply given by the difference between j and i. To learn the distance function, supervised regression can be employed by minimizing the objective.
$\mathcal{L}_{d}=\frac{1}{2}\mathbb{E}_{\tau\sim p_{\tau}}\left[\sum_{i\in[0,T]}\sum_{j\in[i,T]}\left(d_{\psi}(s_{i},s_{j})-(j-i)\right)^{2}\right].$ (12)
To address the GCRL problem, it is crucial to incorporate negative samples to enhance the distance between states and goals that cannot be attained solely through a low-level policy. Additionally, we adopt the objective defined in Equation 4.
C.5 PEBBLE
The key concept of PEBBLE is to learn a reward model on human preference. Unlike the learning strategy of the REWARD model in the MENTOR, it has an unsupervised pretraining step via intrinsic reward. Here, we apply the state entropy $H(s)=-\mathbb{E}\log p(s)$. By converting the policy to a simpler version, the pretraining reward function is set as $\text{log}(||s_{i}-s_{i}^{k}||)$ in the batch. This implies that maximizing the distance between a state and its nearest neighbor increases the overall state entropy.
Figure 10: Experiments on easy domain FetchReach. DDL and PEBBLE successfully learned the task in FetchReach.
Appendix D Further Analysis and Ablations
In this section, we will provide analysis and additional ablations not included in the main body.
D.1 Experiments on FetchReach
[…]
D.2 Heatmap analysis of ablated DDC
The analysis of the heatmaps in Figure 11 shows complex interactions between the distance and reward models in the MENTOR framework. On the right, the oracle reward function in the Four rooms domain indicates where the agent likely earns higher rewards, guiding it from the lower right to the upper right corner. On the left, the distance model assesses subgoal difficulty. The reward heatmap, shaped by human feedback, approximates the oracle function and overlaps with the distance heatmap. This overlap, calculated with $r_{hf}(s,sg,g)+\alpha H(sg,g,k)$, identifies areas optimizing rewards in alignment with the low-level policy’s capabilities. However, without DDC, the human feedback-based reward model is limited, indicating only high-reward areas without guiding the agent on how to reach them.
Figure 11: Evolutions of heatmaps on the Four rooms domain. The color scheme used in the heat map employs red to depict high values and blue to represent low values.
Figure 12: Further ablations
D.3 Ablations of negative sampling in DDC, near policy sampling in RLHF and exploration at the high-level
Tricks in RLHF and distance model. In Figure 12, we evaluate the effect of negative sampling in DDC and near policy sampling in RLHF. The distance heatmap without a negative sample gives an incorrect estimate of the distance from states to some subgoals that the low-level cannot reach. This error may induce the high-level policy to generate subgoals the low-level policy can not complete. The near policy is an additional technique in our framework. By comparing reward heatmaps with and without the inclusion of near-policy samples, we observe that when near-policy samples are absent, the reward model tends to emphasize the entirety of the state space. This, in turn, makes it more challenging and computing resource-consuming for the reward model to effectively learn.
Analysis exploration at the high-level. During the training process of MENTOR, we have observed that there are instances where the high-level policy becomes stuck in local optima as we incrementally increase the value of $k$ for a long time. This phenomenon is attributed to two reasons: 1) It lacks pretraining at the high-level, and 2) the high-level lacks the exploratory ability, leading to not trying new subgoals. Pretraining plays a crucial role in RLHF, yet it can be inefficient at the high-level due to limited samples (often only one per episode). Sampling subgoals across the entire state space becomes computationally wasteful. Instead, the high-level policy does not require knowledge of the optimal subgoal’s location in the entire state space, but only within a specified distance range of $k$. When the high-level policy is optimized within a specified range of $k$, increasing $k$ to $k+\Delta k$ results in the high-level policy getting stuck in local optima. This is because the reward model does not query new sample pairs to provide useful signals. In the operation of the high-levels, we find it necessary to increase the exploratory ability. As shown in Figure 12, reward learning becomes challenging without MaxEnt (maximum entropy). It gets stuck in the learned area and refuses to explore novel regions.
In addition, we compare RND and MaxEnt in terms of Exploratory at high-levels. After conducting experiments displayed in Figure 13, we discovered that MaxEnt outperforms RND. While RND promotes the exploration of novel states, MaxEnt focuses on improving action entropy. The combination of the reward model and DDC can lead to previously disregarded decisions being recognized as good decisions. RND can learn and incorporate these decisions previously, treating them as non-novel at late times. However, MaxEnt does not encounter this issue and as a result, it achieves significantly better performance compared to RND.
Figure 13: Explore encouragement effects analysis on FetchPickAndPlace domain.

### v2

@@ V Conclusion
We recognize the complexity of human guidance beyond just subgoal selection and aim to explore a wider range of feedback integration to enhance learning dynamics. We plan to further expand the framework’s applications and refine its mechanisms, aiming to advance hierarchical reinforcement learning and create more intuitive, adaptable learning systems.
Appendix A Domains Details
@@
$\text{r}_{\text{Four rooms}}(s)=$ (14)
$\begin{cases}-\lVert s-[0.25,0.25]\rVert_{2}&\text{if }s_{x}\geq 0\text{ and }s_{y}\geq 0,\\ -\lVert s-[0,0.3]\rVert_{2}-0.3&\text{if }s_{x}\leq 0\text{ and }s_{y}\geq 0,\\ -\lVert s-[-0.3,0]\rVert_{2}-0.6&\text{if }s_{x}\leq 0\text{ and }s_{y}\leq 0,\\ -\lVert s-[0,-0.3]\rVert_{2}-1&\text{if }s_{x}\geq 0\text{ and }s_{y}\leq 0.\end{cases}$
Appendix B Details of Implementation
Labels acquisition
We obtain labels through two ways: human annotation and script synthesis. The script-synthesized labels are generated based on dense reward functions.
After extracting tuple pairs from the replay buffer $B_{\mathrm{hi}}$ in line 22 of Algorithm 1, such as ($s_{\mathrm{hi}_{1}},g^{\text{sub}}_{1},g_{1}$ and $s_{\mathrm{hi}_{2}},g^{\text{sub}}_{2},g_{2}$), the subsequent step involves obtaining preference labels through interaction with humans or a script. We have designed an interactive interface for human annotation, see Figure 15. This interface is constructed based on the Four rooms domain, where it differentiates tuple pair information through distinct colors of hollow and solid circles. Since the goal position is fixed within this domain, we represent the goal located in the top right corner with a red solid circle. When an annotator needs to determine which tuple is more advantageous for policy learning, they can either click the gray button situated to the right of the letters ”goal” or input ’0’ or ’1’ on the keyboard. If the annotator cannot distinguish which tuple is good or bad, he can enter the ’2’ on the keyboard, and we will set the label to 0.5 to indicate the same preference.
We employ a script to fulfill the role of annotators. The script will calculate the dense reward using an oracle reward function. In the first five domains, we employ the Euclidean distance between the subgoal and the goal. In the Four rooms domain, we use the reward function as mentioned in Equation 14. Depending on the magnitude of the calculated reward for different tuples, we can assign the label to the tuple corresponding to the higher reward.
Figure 15: A simple interface for human annotation.
TABLE III: Hyperparameters of MENTOR.
@@ Appendix B Details of Implementation
Hyperparameters
see Table III.